OpenStudy (anonymous):

Solve for b. 3a + 2b = c Can Someone help me out?

OpenStudy (rushwr):

What did u get ?

OpenStudy (rushwr):

What I got was this: \[b = \frac{ (c- 3a) }{ 2 }\]

OpenStudy (anonymous):

well the choices are b=c-(3a-2) b=(c-3a)/2 b=(3a-c)/2

OpenStudy (anonymous):

and I chose the second one

OpenStudy (rushwr):

Did u get how to do it ?

OpenStudy (rushwr):

correct the answer is the 2nd one :)

OpenStudy (anonymous):

on that one yes, thank you I just wanted to make sure it was right

OpenStudy (rushwr):

Anytym :)